OpenStudy (anonymous):

Some one please solve Layla has allotted $76 for gifts while on vacation. She plans to buy a souvenir for her parents for $16 and spend the rest on 4 T-shirts for her friends. If x represents the price, in dollars, of a T-shirt, which inequality correctly represents the money Layla will spend on souvenirs?

OpenStudy (anonymous):

\[16+4x \le 76\]

OpenStudy (anonymous):

\[x \le 15\]

OpenStudy (anonymous):

thx... the first one was correct... i'll give you a metal

OpenStudy (anonymous):

No worries, I just solved for x for the second one. :)

OpenStudy (anonymous):

If the problem states that Layla will "spend the rest" of her money on those 4 T-shirts, then \[16+4x = 76\] represents the money Layla will spend for gifts. As to answering: "which inequality correctly represents the money Layla will spend on souvenirs?" I would say: \[s = 16\] where s is the money Layla will spend on souvenirs. ...semantics

OpenStudy (anonymous):

Good point Broken Fixer, it depends on how you read the question.